### v1

# Elementary proof of some Ramanujan-type identities

Elementary proof of some Ramanujan-type identities
M.A. Korolev1)1)1)Steklov Mathematical Institute of Russian Academy of Sciences. E-mail: korolevma@mi-ras.ru, hardy_ramanujan@mail.ru
We give an elementary proof of some identities that express the squares of Riemann zeta function at integer points by the series involving hyperbolic functions, digamma function, Bernoulli numbers etc.
Bibliography: 8 titles.
[…]
The main purpose of this note is to apply the general identity established in [1] to derive some relations that connect the squares of Riemann zeta-function $\zeta(z)$ at integer points with digamma-function, Bernoulli numbers etc.
The following lemma is the particular case of general assertion from [1]:
Lemma 1. Suppose that $\{a_{n}\}$, $\{b_{n}\}$, $n=1,2,\ldots,N$, are arbitrary numerical sequences such that $b_{n}>0$ for any $n$. Then the following identity holds:
$\sum\limits_{m,n=1}^{N}a_{m}a_{n}\,\frac{b_{m}}{b_{m}+b_{n}}=\frac{1}{2}\biggl(\sum\limits_{n=1}^{N}a_{n}\biggr)^{2}.$ (1)
For convenience of the reader, we give here the short proof of (1):
$\biggl(\sum\limits_{n=1}^{N}a_{n}\biggr)^{2}=\sum\limits_{m,n=1}^{N}a_{m}a_{n}=\sum\limits_{m,n=1}^{N}a_{m}a_{n}\frac{b_{m}+b_{n}}{b_{m}+b_{n}}=\\ =\sum\limits_{m,n=1}^{N}a_{m}a_{n}\,\frac{b_{m}}{b_{m}+b_{n}}+\sum\limits_{m,n=1}^{N}a_{m}a_{n}\,\frac{b_{n}}{b_{m}+b_{n}}=2\sum\limits_{m,n=1}^{N}a_{m}a_{n}\,\frac{b_{m}}{b_{m}+b_{n}}.\quad\square$
In [1], such type relations were used to establish some amazing identities for the values of Riemann zeta-function. For example, we have
$\sum\limits_{m,n=1}^{+\infty}\frac{1}{m^{2}(m^{2}+n^{2})}=\frac{1}{2}\,\zeta^{2}(2)=\frac{\pi^{4}}{72},\quad\sum\limits_{m,n=1}^{+\infty}\frac{1}{m^{3}(m^{3}+n^{3})}=\frac{1}{2}\,\zeta^{2}(3).$
In what follows, we give some applications of (1) that seem to be new.
Notations. Symbols $k,\ell,m,n,q,r,s$ stand for integers. For $k\geqslant 1$, $r\geqslant 0$ we set
$\varepsilon_{r}=\varepsilon_{r}(k)=\exp{\biggl(\frac{\pi i(2r+1)}{2k}\biggr)},\quad\omega_{r}=\omega_{r}(k)=\exp{\biggl(\frac{\pi i(2r+1)}{2k+1}\biggr)}.$
Next, for real positive $w$ we define
$a_{k}(w)=\frac{1}{k}\sum\limits_{r=0}^{k-1}\pi\varepsilon_{r}\operatorname{ctg}{(\pi\varepsilon_{r}w)},\;\;b_{k}(w)=\frac{1}{k}\sum\limits_{r=0}^{2k-1}\varepsilon_{r}^{2}\psi(\varepsilon_{r}w),\;\;c_{k}(w)=\frac{1}{2k+1}\sum\limits_{r=0}^{2k}\omega_{r}\psi(-\omega_{r}w),$
where $\psi(z)=\Gamma^{\prime}(z)/\Gamma(z)$ is digamma-function. We also use the standard notations for basic arithmetical functions: Möbius function $\mu(n)$, divisor function $\tau(n)$ and its multidimensional analogue $\tau_{\nu}(n)$, Liouville function $\lambda(n)$, von Mangoldt function $\Lambda(n)$. The symbol $\sigma_{k}(n)$ means the sum of $k$th powers of divisors of $n$ and the symbol $\omega(n)$ stands for the number of prime divisors of $n$ counted without multiplicities.
[…]
All the proofs of main theorems are quite elementary. However, they used the following assertion that allows one to change the orders of summation in multiple sums.
Lemma 2. Suppose that $\{\alpha_{m,n}\}$ is a sequence of complex numbers and at least one of the following quantities
$\sum_{m}\sum_{n}|\alpha_{m,n}|,\quad\sum_{n}\sum_{m}|\alpha_{m,n}|$
is finite. Then the double series $\sum\limits_{m,n}\alpha_{m,n}$ converges and its sum equals to
$\sum_{m}\sum_{n}\alpha_{m,n}=\sum_{n}\sum_{m}\alpha_{m,n}.$
[…]
We also need several formulas for the expansions of the simplest rational functions.
Lemma 3. Suppose that $k\geqslant 1$ is integer. Then, for any $s$, $0\leqslant s\leqslant 2k-1$, we have
$\frac{w^{s}}{w^{2k\mathstrut}+1}=\frac{(-1)^{s}}{2k}\sum\limits_{r=0}^{2k-1}\frac{\varepsilon_{r}^{s+1}}{w+\varepsilon_{r}}.$ (2)
Proof. First we prove (2) for $s=0$. To do this, we denote
$P(w)=\frac{1}{2k}\sum\limits_{r=0}^{2k-1}\frac{\varepsilon_{r}}{w-\varepsilon_{r}},\quad Q(w)=(w^{2k}+1)P(w),$
and fix any $\ell$, $0\leqslant\ell\leqslant 2k-1$. Then
$Q(\varepsilon_{\ell})=\lim_{w\to\varepsilon_{\ell}}Q(w)=\frac{\varepsilon_{\ell}}{2k}\lim_{w\to\varepsilon_{t}}\frac{w^{2k}+1}{w-\varepsilon_{\ell}}=\frac{\varepsilon_{\ell}}{2k}\cdot 2k\varepsilon_{\ell}^{2k-1}=\varepsilon_{\ell}^{2k}=-1.$
Thus, the polynomial $Q(w)+1$ has at least $2k$ roots, while its degree does not exceed 2k12k-1. Hence, $Q(w)\equiv-1$ and therefore
$\frac{1}{w^{2k}+1}=-P(w)=-\frac{1}{2k}\sum\limits_{r=0}^{2k-1}\frac{\varepsilon_{r}}{w-\varepsilon_{r}}=\frac{1}{2k}\sum\limits_{r=0}^{2k-1}\frac{\varepsilon_{r}}{w+\varepsilon_{r}}.$
Suppose now that the assertion is proved for any $s$, $0\leqslant s\leqslant n$, where $n\leqslant 2k-2$. Note that for such $n$ we have
$\sum\limits_{r=0}^{2k-1}\varepsilon_{r}^{n+1}=\exp{\biggl(\frac{\pi i(n+1)}{2k}\biggr)}\sum\limits_{r=0}^{2k-1}\exp{\biggl(\frac{2\pi i(n+1)r}{2k}\biggr)}=0.$
Hence,
$\frac{(-1)^{n+1}}{2k}\sum\limits_{r=0}^{2k-1}\frac{\varepsilon_{r}^{n+1}}{w}=0.$
Adding this relation term by term to the identity
$\frac{(-1)^{n}}{2k}\sum\limits_{r=0}^{2k-1}\frac{\varepsilon_{r}^{n+1}}{w+\varepsilon_{r}}=\frac{w^{n}}{w^{2k\mathstrut}+1},$
we get
$\frac{w^{n}}{w^{2k}+1}=\frac{(-1)^{n+1}}{2k}\sum\limits_{r=0}^{2k-1}\varepsilon_{r}^{n+1}\biggl(\frac{1}{w}-\frac{1}{w+\varepsilon_{r}}\biggr)=\frac{(-1)^{n+1}}{2k}\sum\limits_{r=0}^{2k-1}\frac{\varepsilon_{r}^{n+2}}{w(w+\varepsilon_{r})}.$
Multiplying both parts of the last equality by $w$, we arrive at the desired assertion. $\square$
Corollary 1. Under assumptions of Lemma 3, for any $m,n\geqslant 1$ we have
$\frac{n^{2k-1}}{m^{2k\mathstrut}+n^{2k\mathstrut}}=\frac{1}{2k}\sum\limits_{r=0}^{k-1}\varepsilon_{r}\biggl(\frac{1}{m+n\varepsilon_{r}}-\frac{1}{m-n\varepsilon_{r}}\biggr).$
Proof. Taking $s=1$, $w=m/n$ in (2), we get
$\frac{n^{2k-1}}{m^{2k\mathstrut}+n^{2k\mathstrut}}=\frac{1}{2k}\sum\limits_{r=0}^{2k-1}\frac{\varepsilon_{r}}{m+n\varepsilon_{r}}=\frac{1}{2k}\sum\limits_{r=0}^{k-1}\biggl(\frac{\varepsilon_{r}}{m+n\varepsilon_{r}}+\frac{\varepsilon_{r+k}}{m+n\varepsilon_{r+k}}\biggr).$
Now it suffices to note that $\varepsilon_{r+k}=-\varepsilon_{r}$. $\square$
Corollary 2. Under the assumptions of Lemma 3, for any $m,n\geqslant 1$ and for any $k,\ell$, $0\leqslant\ell\leqslant 2k-1$, we have
$\frac{m^{\ell}n^{2k-1-\ell}}{m^{2k\mathstrut}+n^{2k\mathstrut}}=\frac{(-1)^{\ell+1}}{2k}\sum\limits_{r=0}^{2k-1}\varepsilon_{r}^{\ell+1}\biggl(\frac{1}{m}-\frac{1}{m+n\varepsilon_{r}}\biggr).$
Proof. Taking $s=\ell$, $w=m/n$ in (2), we get
$\frac{m^{\ell}n^{2k-1-\ell}}{m^{2k\mathstrut}+n^{2k\mathstrut}}=\frac{(-1)^{\ell}}{2k}\sum\limits_{r=0}^{2k-1}\frac{\varepsilon_{r}^{\ell+1}}{m+n\varepsilon_{r}}.$
Subtracting this identity from the equality
$\frac{(-1)^{\ell+1}}{2k}\sum\limits_{r=0}^{2k-1}\frac{\varepsilon_{r}^{\ell+1}}{m}=0,$
we easily obtain the desired assertion. $\square$
Lemma 4. For any $k\geqslant 1$ and $s$, $0\leqslant s\leqslant 2k$, one has
$\frac{w^{s}}{w^{2k+1}+1}=-\frac{1}{2k+1}\sum\limits_{r=0}^{2k}\frac{\omega_{r}^{s+1}}{w-\omega_{r}}.$
Corollary 1. Under the assumptions of Lemma 4, for any $m,n\geqslant 1$ one has
$\frac{n^{2k}}{m^{2k+1\mathstrut}+n^{2k+1\mathstrut}}=\frac{1}{2k+1}\sum\limits_{r=0}^{2k}\omega_{r}\biggl(\frac{1}{m}-\frac{1}{m-n\omega_{r}}\biggr).$
Corollary 2. Under the assumptions of Lemma 4, for any $m,n\geqslant 1$ and for any $k\geqslant 1$, $\ell$, $0\leqslant\ell\leqslant 2k$, one has
$\frac{m^{\ell}n^{2k-\ell}}{m^{2k+1\mathstrut}+n^{2k+1\mathstrut}}=-\,\frac{1}{2k+1}\sum\limits_{r=0}^{2k}\frac{\omega_{r}^{\ell}}{m-n\omega_{r}}.$
Proofs of these relations follow word-by-word to the proofs of Lemma 3 and its corollaries.
In what follows, we will use the basic properties of digamma function $\psi(z)$.
Lemma 5. Let $z$ be a complex number such that $z\neq-m$, $m=0,1,2,3,\ldots.$ Then
$\psi(z)=-\gamma-\frac{1}{z}+\sum\limits_{n=1}^{+\infty}\biggl(\frac{1}{n}-\frac{1}{n+z}\biggr).$
Lemma 6. For any non-integer complex $z$, one has
$\psi(-z)=\psi(z)+\frac{1}{z}+\pi\operatorname{ctg}{(\pi z)}.$
Lemma 7. For any complex non-integer $z$, one has
$\sum\limits_{n=1}^{+\infty}\biggl(\frac{1}{n+z}-\frac{1}{n-z}\biggr)=\pi\operatorname{ctg}{(\pi z)}-\frac{1}{z}.$
[…]
Lemma 8. Suppose that $z=\varrho e^{i\varphi}$, where $\varrho>0$ and $0<|\varphi|<\pi$ are real numbers. Then the following inequality holds:
$\biggl|\operatorname{ctg}{(\pi z)}-\frac{1}{\pi z}\biggr|<\frac{3}{2}\cdot\frac{\min{\bigl(1,\varrho\bigr)}}{|\!\sin{\varphi}|}.$
Proof. By Lemma 7,
$\biggl|\operatorname{ctg}{(\pi z)}-\frac{1}{\pi z}\biggr|\leqslant\frac{1}{\pi}\sum\limits_{n=1}^{+\infty}\biggl|\frac{1}{n+z}-\frac{1}{n-z}\biggr|=\frac{2|z|}{\pi}\sum\limits_{n=1}^{+\infty}\frac{1}{|n^{2}-z^{2}|}.$
Further,
$\bigl|n^{2}-z^{2}\bigr|^{2}=\bigl|n^{2}-\varrho^{2}e^{2i\varphi}\bigr|^{2}=n^{4}+\varrho^{4}-2(n\varrho)^{2}\cos{(2\varphi)}.$
If $\cos{(2\varphi)}\leqslant 0$ then
$\bigl|n^{2}-z^{2}\bigr|^{2}\geqslant n^{4}+\varrho^{4}\geqslant\frac{1}{2}\,(n^{2}+\varrho^{2})^{2},$
so we have
$\bigl|n^{2}-z^{2}\bigr|\geqslant\frac{n^{2}+\varrho^{2}}{\sqrt{2}}.$ (3)
Suppose now that $\cos{(2\varphi)}>0$. Then
$\bigl|n^{2}-z^{2}\bigr|^{2}\geqslant n^{4}+\varrho^{4}-(n^{4}+\varrho^{4})\cos{(2\varphi)}=(n^{4}+\varrho^{4})\cdot 2\sin^{2}{\varphi}\geqslant(n^{2}+\varrho^{2})^{2}\sin^{2}{\varphi},$
and therefore
$\bigl|n^{2}-z^{2}\bigr|\geqslant|\!\sin{\varphi}|(n^{2}+\varrho^{2}).$ (4)
From (3) and (4) it follows that in both cases we have
$\bigl|n^{2}-z^{2}\bigr|\geqslant\frac{|\!\sin{\varphi}|}{\sqrt{2}}\,(n^{2}+\varrho^{2})$
Hence,
$\biggl|\operatorname{ctg}{(\pi z)}-\frac{1}{\pi z}\biggr|\leqslant\frac{2\sqrt{2}}{\pi|\!\sin{\varphi}|}\sum\limits_{n=1}^{+\infty}\frac{\varrho}{n^{2}+\varrho^{2}}.$
In the case $0<\varrho<3/\pi$, the sum over $n$ is less than
$\varrho\sum\limits_{n=1}^{+\infty}\frac{1}{n^{2}}=\frac{\pi^{2}\varrho}{6}=\frac{\pi}{2}\cdot\frac{\pi\varrho}{3}=\frac{\pi}{2}\,\min{\biggl(1,\frac{\pi\varrho}{3}\biggr)}.$
If $\varrho\geqslant 3/\pi$ then this sum does not exceed
$\int_{0}^{+\infty}\frac{\varrho\,dx}{x^{2}+\varrho^{2}}=\frac{\pi}{2}=\frac{\pi}{2}\,\min{\biggl(1,\frac{\pi\varrho}{3}\biggr)}.$
Finally we get:
$\biggl|\operatorname{ctg}{(\pi z)}-\frac{1}{\pi z}\biggr|\leqslant\frac{2\sqrt{2}}{\pi|\!\sin{\varphi}|}\cdot\frac{\pi}{2}\,\min{\biggl(1,\frac{\pi\varrho}{3}\biggr)}\leqslant\frac{\pi\sqrt{2}}{3}\cdot\frac{\min{\bigl(1,\varrho\bigr)}}{|\!\sin{\varphi}|}<\frac{3}{2}\cdot\frac{\min{\bigl(1,\varrho\bigr)}}{|\!\sin{\varphi}|}.$
Lemma is proved. $\square$
Lemma 9. For any complex $z$ such that $|\!\arg{z}|<\pi/2$ and for any integer $M\geqslant 1$ the following identity holds:
$\psi(z)=\ln{z}-\frac{1}{2z}-\sum\limits_{m=1}^{M}\frac{B_{2m}}{2m}\cdot\frac{1}{z^{2m\mathstrut}}+\frac{2(-1)^{M+1}}{z^{2M\mathstrut}}\,j_{M}(z).$
Here
$j_{M}(z)=\int_{0}^{+\infty}\frac{t^{2M+1}}{e^{2\pi t}-1}\,\frac{dt}{(t^{2}+z^{2})}$
and $B_{j}$ are Bernoulli numbers: $B_{2}=1/2$, $B_{4}=-1/30$, $B_{6}=1/42$, $\ldots$, $B_{2n+1}=0$ for $n\geqslant 1$.
[…]
Lemma 10. For any complex non-integer $z=x+iy$, one has
$\operatorname{ctg}{(x+iy)}=\frac{\sin{(2x)}-i\operatorname{sh}{(2y)}}{2(\operatorname{sh}^{2}{(y)}+\sin^{2}{(x)})}=\frac{\sin{(2x)}-i\operatorname{sh}{(2y)}}{\operatorname{ch}{(2y)}-\cos{(2x)}}.$
[…]
Here we prove some identities that express the squares of Riemann zeta function at integer points by the sums of some rapidly convergent series.
Theorem 1. For any integer $k\geqslant 1$, the following relation holds:
$\zeta^{2}(2k)+\zeta(4k)=\sum\limits_{n=1}^{+\infty}\frac{a_{k}(n)}{n^{4k-1\mathstrut}}.$
Proof. Fixing any $m,n\geqslant 1$ and using the Corollary 1 of Lemma 3, we get
$\frac{n^{2k-1}}{m^{2k}+n^{2k}}=\frac{1}{2k}\sum\limits_{r=0}^{k-1}\varepsilon_{r}\biggl(\frac{1}{m+n\varepsilon_{r}}-\frac{1}{m-n\varepsilon_{r}}\biggr).$
Hence,
$\sum\limits_{m=1}^{+\infty}\frac{n^{2k-1}}{m^{2k}+n^{2k}}=\frac{1}{2k}\sum\limits_{m=1}^{+\infty}\sum\limits_{r=0}^{k-1}\varepsilon_{r}\biggl(\frac{1}{m+n\varepsilon_{r}}-\frac{1}{m-n\varepsilon_{r}}\biggr).$ (5)
Further, (4) implies that
$\bigl|m^{2}-(n\varepsilon_{r})^{2}\bigr|\geqslant(m^{2}+n^{2})\sin{\frac{\pi}{2k}}.$
Hence,
$\biggl|\varepsilon_{r}\biggl(\frac{1}{m+n\varepsilon_{r}}-\frac{1}{m-n\varepsilon_{r}}\biggr)\biggr|=\frac{2n}{|m^{2\mathstrut}-(n\varepsilon_{r})^{2\mathstrut}|}\leqslant\biggl(\sin{\frac{\pi}{2k\mathstrut}}\biggr)^{-1}\cdot\frac{2n}{m^{2}+n^{2\mathstrut}},.$ (6)
Lemma 2 together with (6) allow one to change the order of summation in the right-hand side of (5). Using the identity of Lemma 7, we find
$\sum\limits_{m=1}^{+\infty}\frac{n^{2k-1}}{m^{2k\mathstrut}+n^{2k\mathstrut}}=\frac{1}{2k}\sum\limits_{r=0}^{k-1}\varepsilon_{r}\sum\limits_{m=1}^{+\infty}\biggl(\frac{1}{m+n\varepsilon_{r}}-\frac{1}{m-n\varepsilon_{r}}\biggr)=\\ =\frac{1}{2k}\sum\limits_{r=0}^{k-1}\varepsilon_{r}\biggl(\pi\operatorname{ctg}{(\pi\varepsilon_{r}n)}-\frac{1}{n\varepsilon_{r}}\biggr)=\frac{1}{2}\,a_{k}(n)-\frac{1}{2n}.$ (7)
Further, by definition of $a_{k}(n)$ and Lemma 8 we have
$\biggl|a_{k}(n)-\frac{1}{n}\biggr|=\biggl|\frac{\pi}{k}\sum\limits_{r=0}^{k-1}\varepsilon_{r}\biggl(\operatorname{ctg}{(\pi\varepsilon_{r}n)}-\frac{1}{\pi n\varepsilon_{r}}\biggr)\biggr|\leqslant\frac{\pi}{k}\sum\limits_{r=0}^{k-1}\biggl|\operatorname{ctg}{(\pi\varepsilon_{r}n)}-\frac{1}{\pi n\varepsilon_{r}}\biggr|\ll_{k}1,$
$\bigl|a_{k}(n)\bigr|\ll_{k}1.$
Multiplying both parts of (7) by $n^{-(4k-1)}$ and summing over $n\geqslant 1$, we obtain
$\sum\limits_{m,n=1}^{+\infty}\frac{1}{n^{2k}(m^{2k\mathstrut}+n^{2k\mathstrut})}=\frac{1}{2}\sum\limits_{n=1}^{+\infty}\frac{a_{k}(n)}{n^{4k-1}}-\frac{1}{2}\,\zeta(4k).$
By Lemma 1, left hand side coincides with
$\sum\limits_{m,n=1}^{+\infty}\frac{1}{(mn)^{2k}}\cdot\frac{m^{2k}}{m^{2k\mathstrut}+n^{2k\mathstrut}}=\frac{1}{2}\biggl(\,\sum\limits_{m=1}^{+\infty}\frac{1}{m^{2k\mathstrut}}\biggr)^{2}=\frac{1}{2}\,\zeta^{2}(2k).$
Thus Theorem 1 follows. $\square$
Corollary. For any $k\geqslant 1$, the following identity holds:
$\zeta^{2}(2k)+\zeta(4k)=\frac{\pi}{k}\,\frac{\zeta(4k-1)}{s(0)}+\frac{\pi}{2k}\sum\limits_{n=1}^{+\infty}\frac{\alpha_{k}(n)}{n^{4k-1\mathstrut}}.$ (8)
Here
$\alpha_{k}(n)=\sum\limits_{r=0}^{k-1}\frac{c(r)\sin{(2\pi nc(r))}+s(r)\cos{(2\pi nc(r))}-s(r)e^{-2\pi ns(r)}}{\operatorname{sh}^{2}{(\pi ns(r))}+\sin^{2}{(\pi nc(r))}},\\ c(r)=\cos\biggl(\frac{\pi}{2k}(2r+1)\biggr),\quad s(r)=\sin\biggl(\frac{\pi}{2k}(2r+1)\biggr).$
Remark. Setting $k=1$ in (8) and using the relations $\zeta(2)=\pi^{2}/6$, $\zeta(4)=\pi^{4}/90$, after obvious transformations, we get <<Cauchy-Lerch-Ramanujan formula>>2)2)2)This name is not generally accepted.
$\zeta(3)=\frac{7\pi^{3}}{180}-2\sum\limits_{n=1}^{+\infty}\frac{1}{n^{3\mathstrut}(e^{2\pi n}-1)}$
(see [2, formula (120)], [3] and [4, pp. 275-276, 293]).
Theorem 2. For any $k,\ell\geqslant 1$ such that $1\leqslant\ell\leqslant 2k-2$, the following relation holds:
$\zeta^{2}(2k-\ell)=\sum\limits_{n=1}^{+\infty}\frac{b_{k,\ell}(n)}{n^{4k-2\ell-1}},\quad\text{where}\quad b_{k,\ell}(n)=\frac{(-1)^{\ell+1}}{k}\sum\limits_{r=0}^{2k-1}\varepsilon_{r}^{\ell+1}\psi(n\varepsilon_{r}).$
Proof. By Corollary 2 of Lemma 3, for any $m,n\geqslant 1$ we have
$\frac{m^{\ell}n^{2k-\ell-1}}{m^{2k\mathstrut}+n^{2k\mathstrut}}=\frac{(-1)^{\ell+1}}{2k}\sum\limits_{r=0}^{2k-1}\varepsilon_{r}^{\ell+1}\biggl(\frac{1}{m}-\frac{1}{m+n\varepsilon_{r}}\biggr).$
Summing both part over $m$ and using Lemma 5, we get
$\sum\limits_{m=1}^{+\infty}\frac{m^{\ell}n^{2k-\ell-1}}{m^{2k\mathstrut}+n^{2k\mathstrut}}=\frac{(-1)^{\ell+1}}{2k}\sum\limits_{m=1}^{+\infty}\sum\limits_{r=0}^{2k-1}\varepsilon_{r}^{\ell+1}\biggl(\frac{1}{m}-\frac{1}{m+n\varepsilon_{r}}\biggr)=\\ =\frac{(-1)^{\ell+1}}{2k}\sum\limits_{r=0}^{2k-1}\varepsilon_{r}^{\ell+1}\sum\limits_{m=1}^{+\infty}\biggl(\frac{1}{m}-\frac{1}{m+n\varepsilon_{r}}\biggr)=\frac{(-1)^{\ell+1}}{2k}\sum\limits_{r=0}^{2k-1}\varepsilon_{r}^{\ell+1}\biggl(\psi(n\varepsilon_{r})+\gamma+\frac{1}{n\varepsilon_{r}}\biggr)=\\ =\frac{(-1)^{\ell+1}}{2k}\sum\limits_{r=0}^{2k-1}\varepsilon_{r}^{\ell+1}\psi(n\varepsilon_{r})=\frac{1}{2}\,b_{k,\ell}(n)$ (9)
(the change of orders of summation is easily justified as above).
Further, we multiply both parts of (9) by $n^{-(4k-2\ell-1)}$ and sum over $n$. Thus, using Lemma 1, we get
$\frac{1}{2}\sum\limits_{n=1}^{+\infty}\frac{b_{k,\ell}(n)}{n^{4k-2\ell-1}}=\sum\limits_{m,n=1}^{+\infty}\frac{1}{n^{4k-2\ell-1}}\cdot\frac{m^{\ell}n^{2k-\ell-1}}{m^{2k\mathstrut}+n^{2k\mathstrut}}=\sum\limits_{m,n=1}^{+\infty}\frac{m^{\ell}}{n^{2k-\ell}(m^{2k\mathstrut}+n^{2k\mathstrut})}=\\ =\sum\limits_{m,n=1}^{+\infty}\frac{1}{(mn)^{2k-\ell}}\cdot\frac{m^{2k}}{(m^{2k\mathstrut}+n^{2k\mathstrut})}=\frac{1}{2}\biggl(\,\sum\limits_{m=1}^{+\infty}\frac{1}{m^{2k-\ell}}\biggr)^{2}=\frac{1}{2}\,\zeta^{2}(2k-\ell).$
Theorem is proved. $\square$
Corollary 1. The following identity holds:
$\zeta^{2}(3)=-\sum\limits_{n=1}^{+\infty}\frac{\beta(n)}{n^{5}},\quad\beta(n)=\operatorname{Im}{\bigl(\psi(n\varepsilon_{0})+\psi(-n\varepsilon_{0})\bigr)},\quad\varepsilon_{0}=e^{\pi i/4}.$
Proof. Setting $k=2$, $\ell=1$, we get
$\varepsilon_{0}=e^{\pi i/4},\quad\varepsilon_{1}=-\overline{\varepsilon}_{0},\quad\varepsilon_{2}=-\varepsilon_{0},\quad\varepsilon_{3}=\overline{\varepsilon}_{0}.$
Further, we have $\varepsilon_{r}^{2}=(-1)^{r}\cdot i$ for $r=0,1,2$. Hence,
$b_{2,1}(n)=\frac{1}{2}\sum\limits_{r=0}^{3}\varepsilon_{r}^{2}\psi(n\varepsilon_{r})=\frac{i}{2}\sum\limits_{r=0}^{3}(-1)^{r}\psi(n\varepsilon_{r})=\\ =\frac{i}{2}\bigl(\psi(n\varepsilon_{0})-\psi(n\varepsilon_{1})+\psi(n\varepsilon_{2})-\psi(n\varepsilon_{3})\bigr)=\frac{i}{2}\bigl(\psi(n\varepsilon_{0})-\psi(-n\overline{\varepsilon}_{0})+\psi(-n\varepsilon_{0})-\psi(n\overline{\varepsilon}_{0})\bigr)=\\ =-\operatorname{Im}{\bigl(\psi(n\varepsilon_{0})+\psi(-n\varepsilon_{0})\bigr)}.\quad\square$
Corollary 2. For any fixed $m\geqslant 0$, the following identity holds:
$\zeta^{2}(3)=\frac{\pi}{2}\,\zeta(5)-\sum\limits_{r=0}^{m}\frac{(-1)^{r}B_{4r+2}}{2r+1}\,\zeta(4r+7)+\mathfrak{S}_{0}+(-1)^{m}\int_{0}^{+\infty}\frac{G_{m}(t)dt}{e^{2\pi t}-1}.$
Here
$\mathfrak{S}_{0}=\pi\sum\limits_{n=1}^{+\infty}\frac{1}{n^{5}}\cdot\frac{\cos{(\pi n\sqrt{2})}-e^{-\pi n\sqrt{2}}}{\operatorname{ch}{(\pi n\sqrt{2})}-\cos{(\pi n\sqrt{2})}}=-0.0204388172\ldots,\quad G_{m}(t)=4\sum\limits_{n=1}^{\infty}\frac{(t/n)^{4m+5}}{n^{2\mathstrut}(n^{4\mathstrut}+t^{4\mathstrut})}.$
Proof. By Lemma 6,
$\psi(n\varepsilon_{0})+\psi(-n\varepsilon_{0})=2\psi(n\varepsilon_{0})+\frac{1}{n\varepsilon_{0}}+\pi\operatorname{ctg}{(\pi\varepsilon_{0}n)}.$
Further, Lemma 10 implies that
$\operatorname{ctg}{(\pi\varepsilon_{0}n)}=\frac{\sin{(\pi n\sqrt{2})}-i\operatorname{sh}{(\pi n\sqrt{2})}}{\operatorname{ch}{(\pi n\sqrt{2})}-\cos{(\pi n\sqrt{2})}}.$
Hence,
$\beta(n)=2\operatorname{Im}{\psi(n\varepsilon_{0})}-\frac{1}{n\sqrt{2}}-\frac{\pi\operatorname{sh}{(\pi n\sqrt{2})}}{\operatorname{ch}{(\pi n\sqrt{2})}-\cos{(\pi n\sqrt{2})}}=\\ =2\operatorname{Im}{\psi(n\varepsilon_{0})}-\frac{1}{n\sqrt{2}}-\pi\biggl(1+\frac{\cos{(\pi n\sqrt{2})}-e^{-\pi n\sqrt{2}}}{\operatorname{ch}{(\pi n\sqrt{2})}-\cos{(\pi n\sqrt{2})}}\biggr).$
Taking $z=n\varepsilon_{0}$, $M=2m+1$ in Lemma 9, we find
$\psi(n\varepsilon_{0})=\ln{n}+\frac{\pi i}{4}+\frac{i-1}{2n\sqrt{2}}-\sum\limits_{s=1}^{2m+1}\frac{B_{2s}}{2s}\cdot\frac{i^{-s}}{n^{2s}}+(-1)^{m+1}\cdot\frac{2i}{n^{4m+2\mathstrut}}\int_{0}^{+\infty}\frac{t^{4m+3}}{e^{2\pi t}-1}\cdot\frac{dt}{t^{2\mathstrut}+in^{2\mathstrut}}$
and therefore
$\operatorname{Im}\psi(n\varepsilon_{0})=\frac{\pi}{4}+\frac{1}{2n\sqrt{2}}-\sum\limits_{s=1}^{2m+1}\frac{B_{2s}}{2s}\cdot\frac{\operatorname{Im}{(i^{-s})}}{n^{2s}}+(-1)^{m+1}\cdot\frac{2}{n^{4m+2\mathstrut}}\int_{0}^{+\infty}\frac{t^{4m+5}}{e^{2\pi t}-1}\cdot\frac{dt}{t^{4\mathstrut}+n^{4\mathstrut}}.$
The non-zero contribution to the sum over $s$ is given by odd $s=2r+1$ only. For such $s$, we have $\operatorname{Im}{(i^{-s})}=(-1)^{r+1}$. Thus we obtain:
$\beta(n)=-\frac{\pi}{2}+\sum\limits_{r=0}^{m}\frac{(-1)^{r}B_{4r+2}}{2r+1}\cdot\frac{1}{n^{4r+2}}-\pi\,\frac{\cos{(\pi n\sqrt{2})}-e^{-\pi n\sqrt{2}}}{\operatorname{ch}{(\pi n\sqrt{2})}-\cos{(\pi n\sqrt{2})}}+\\ +\frac{4(-1)^{m+1}}{n^{4m+2\mathstrut}}\int_{0}^{+\infty}\frac{t^{4m+5}}{e^{2\pi t}-1}\cdot\frac{dt}{t^{4\mathstrut}+n^{4\mathstrut}}.$
Now the assertion follows. $\square$
Theorem 3. For any $k\geqslant 1$ the following identity holds:
$\frac{1}{2}\,\zeta^{2}(2k+1)+\zeta(4k+2)=\sum\limits_{n=1}^{+\infty}\frac{c_{k}(n)}{n^{4k+1}}.$
Proof. By Corollary 1 of Lemma 4, for any $m,n\geqslant 1$ one has:
$\frac{n^{2k}}{n^{2k+1\mathstrut}+m^{2k+1\mathstrut}}=\frac{1}{2k+1}\sum\limits_{r=0}^{2k}\omega_{r}\biggl(\frac{1}{m}-\frac{1}{m-n\omega_{r}}\biggr).$ (10)
Summing both parts of (10) over $m\geqslant 1$ and using Lemma 5 we find
$\sum\limits_{m=1}^{+\infty}\frac{n^{2k}}{n^{2k+1\mathstrut}+m^{2k+1\mathstrut}}=\frac{1}{2k+1}\sum\limits_{r=0}^{2k}\omega_{r}\biggl(\psi(-n\omega_{r})+\gamma-\frac{1}{n\omega_{r}}\biggr)=\\ =\frac{1}{2k+1}\sum\limits_{r=0}^{2k}\omega_{r}\psi(-n\omega_{r})-\frac{1}{n}=c_{k}(n)-\frac{1}{n}.$ (11)
The change of orders of summation is easily justified as above, but now we have to use the inequality
$|m-n\omega_{r}|\geqslant(m+n)\sin{\frac{\pi}{2(2k+1)}}.$
Further,
$\biggl|c_{k}(n)-\frac{1}{n}\biggr|\leqslant\frac{1}{2k+1}\sum\limits_{r=0}^{2k}\biggl|\psi(-n\omega_{r})+\gamma-\frac{1}{n\omega_{r}}\biggr|=\frac{1}{2k+1}\sum\limits_{r=0}^{2k}\biggl|\sum\limits_{m=1}^{+\infty}\biggl(\frac{1}{m}-\frac{1}{m-n\varepsilon_{r}}\biggr)\biggr|\leqslant\\ \leqslant\frac{1}{2k+1}\sum\limits_{r=0}^{2k}\sum\limits_{m=1}^{+\infty}\frac{n}{m|m-n\omega_{r}|}\leqslant\biggl((2k+1)\sin{\frac{\pi}{2(2k+1)}}\biggr)^{-1}\sum\limits_{r=0}^{2k}\sum\limits_{m=1}^{+\infty}\frac{n}{m(m+n)}=\\ =\biggl(\sin{\frac{\pi}{2(2k+1)}}\biggr)^{-1}\sum\limits_{m=1}^{+\infty}\biggl(\frac{1}{m}-\frac{1}{m+n}\biggr)=\biggl(\sin{\frac{\pi}{2(2k+1)}}\biggr)^{-1}H_{n},$
where
$H_{n}=1+\frac{1}{2}+\frac{1}{3}+\ldots+\frac{1}{n}\leqslant\ln{n}+1.$
Hence,
$|c_{k}(n)|\leqslant\ln{n}+1+\frac{1}{n}\leqslant\ln{n}+2.$ (12)
Multiplying both parts of (11) by $n^{-(4k+1)}$ and summing over $n$, we get
$\sum\limits_{n=1}^{+\infty}\frac{1}{n^{4k+1}}\biggl(c_{k}(n)-\frac{1}{n}\biggr)=\sum\limits_{m,n=1}^{+\infty}\frac{1}{n^{4k+1}}\cdot\frac{n^{2k}}{m^{2k+1\mathstrut}+n^{2k+1\mathstrut}}.$ (13)
In view of the estimate (12), the left-hand side of (13) equals to
$\sum\limits_{n=1}^{+\infty}\frac{c_{k}(n)}{n^{4k+1\mathstrut}}-\zeta(4k+2).$
By Lemma 1, the right-hand side of (13) coincides with
$\sum\limits_{m,n=1}^{+\infty}\frac{m^{2k+1}}{(mn)^{2k+1\mathstrut}(m^{2k+1\mathstrut}+n^{2k+1\mathstrut})}=\frac{1}{2}\biggl(\,\sum\limits_{m=1}^{+\infty}\frac{1}{m^{2k+1\mathstrut}}\biggr)^{2}=\frac{1}{2}\,\zeta^{2}(2k+1).$
Now theorem follows. $\square$
Corollary 1. The following identity holds:
$\zeta^{2}(3)=\frac{2\pi}{\sqrt{3}}\,\zeta(5)-\frac{2}{3}\,\zeta(6)+\mathfrak{S}+2\sum\limits_{n=1}^{+\infty}\frac{\alpha(n)}{n^{5}}.$
Here
[…]
$\alpha(n)=\frac{2}{3}\,\operatorname{Re}{\bigl(\omega_{0}\psi(n\omega_{0})\bigr)}-\frac{1}{3}\,\psi(n),\quad\omega_{0}=e^{\pi i/3}.$
Proof. Taking $k=1$ we get
$\omega_{r}=e^{\pi i(2r+1)/3}\quad\text{and therefore}\quad\omega_{0}=e^{\pi i/3},\quad\omega_{1}=-1,\quad\omega_{2}=\overline{\omega}_{0}=e^{-\pi i/3}.$
Hence,
$c_{1}(n)=\frac{1}{3}\sum\limits_{r=0}^{2}\omega_{r}\psi(-n\omega_{r})=\frac{2}{3}\,\operatorname{Re}{\bigl(\omega_{0}\psi(-n\omega_{0})\bigr)}-\frac{1}{3}\,\psi(n).$
By Lemma 6,
$\omega_{0}\psi(-n\omega_{0})=\omega_{0}\psi(n\omega_{0})+\frac{1}{n}+\pi\omega_{0}\operatorname{ctg}{(\pi\omega_{0}n)},$
so we have
$c_{1}(n)=\alpha(n)+\frac{2}{3n}+\frac{2}{3}\,\operatorname{Re}{\bigl(\pi\omega_{0}\operatorname{ctg}{(\pi\omega_{0}n)}\bigr)}.$
Next, Lemma 10 implies that
$\operatorname{ctg}{(\pi\omega_{0}n)}=\operatorname{ctg}{\biggl(\frac{\pi n}{2}+\frac{\pi in}{2}\sqrt{3}\biggr)}=-\,\frac{i\operatorname{sh}{(\pi n\sqrt{3})}}{\operatorname{ch}(\pi n\sqrt{3})-\cos(\pi n)}$
and therefore
$\operatorname{Re}{\bigl(\pi\omega_{0}\operatorname{ctg}{(\pi\omega_{0}n)}\bigr)}=\frac{\pi\sqrt{3}}{2}\cdot\frac{\operatorname{sh}{(\pi n\sqrt{3})}}{\operatorname{ch}(\pi n\sqrt{3})-\cos(\pi n)}=\\ =\frac{\pi\sqrt{3}}{2}\cdot\biggl(1+\frac{\cos{(\pi n)}-e^{-\pi n\sqrt{3}}}{\operatorname{ch}(\pi n\sqrt{3})-\cos(\pi n)}\biggr).$
Now it is not difficult to check that the second term in brackets coincides with
$\frac{(-1)^{n}e^{-\pi n\sqrt{3}/2}}{\varphi_{n}\bigl(\pi n\sqrt{3}/2\bigr)}.$
Thus,
$c_{1}(n)=\alpha(n)+\frac{2}{3n}+\frac{\pi}{\sqrt{3}}\biggl(1+\frac{(-1)^{n}e^{-\pi n\sqrt{3}/2}}{\varphi_{n}\bigl(\pi n\sqrt{3}/2\bigr)}\biggr).$
Thus the desired assertion follows. $\square$
Corollary 2. For any fixed $m\geqslant 0$, the following identity holds:
$\zeta^{2}(3)+\zeta(6)=\frac{4\pi}{3\sqrt{3}}\,\zeta(5)+\sum\limits_{r=0}^{m}\frac{B_{6r+4}}{3r+2}\,\zeta(6r+9)+\mathfrak{S}+(-1)^{m}\int_{0}^{+\infty}\frac{F_{m}(t)dt}{e^{2\pi t}-1},$
where $\mathfrak{S}$ is defined in Corollary 1 and
$F_{m}(t)=4\sum\limits_{n=1}^{\infty}\frac{(t/n)^{6m+9}}{n^{6\mathstrut}+t^{6\mathstrut}}.$
Proof. Taking $M=3m+2$ in Lemma 9, after easy calculations we get
$\operatorname{Re}{\bigl(\omega_{0}\psi(n\omega_{0})\bigr)}=\frac{1}{2}\,\ln{n}-\frac{\pi}{2\sqrt{3}}-\frac{1}{2n}-\sum\limits_{s=1}^{3m+2}\frac{B_{2s}}{2s}\cdot\frac{\operatorname{Re}{\bigl(\omega_{0}^{1-2s}\bigr)}}{n^{2s\mathstrut}}+\\ +\frac{2(-1)^{m}}{n^{6m+4}}\int_{0}^{+\infty}\operatorname{Re}{\biggl(\frac{1}{t^{2}+n^{2}\omega_{0}^{2}}\biggr)}\frac{t^{6m+5}\,dt}{e^{2\pi t\mathstrut}-1},\\ \psi(n)=\ln{n}-\frac{1}{2n}-\sum\limits_{s=1}^{3m+2}\frac{B_{2s}}{2s}\cdot\frac{1}{n^{2s}}+\frac{2(-1)^{m}}{n^{6m+4}}\int_{0}^{+\infty}\frac{t^{6m+5}}{e^{2\pi t\mathstrut}-1}\,\frac{dt}{t^{2}+n^{2}}.$
Thus we find that
$\alpha(n)=-\,\frac{\pi}{3\sqrt{3}}-\frac{1}{6n}+\frac{1}{6}\sum\limits_{s=1}^{3m+2}\frac{B_{2s}}{sn^{2s}}\bigl(1-2\operatorname{Re}{(\omega_{0}^{1-2s})}\bigr)+\frac{2}{3}\frac{(-1)^{m}}{n^{6m+4}}\int_{0}^{+\infty}\frac{t^{6m+5}}{e^{2\pi t\mathstrut}-1}\,g_{n}(t)dt,$
where
$g_{n}(t)=\frac{1}{t^{2}+n^{2}}+2\operatorname{Re}{\biggl(\frac{1}{t^{2}+n^{2}\omega_{0}^{2}}\biggr)}$
Note that
[…]
Hence, the non-zero contribution to the sum over $s$ is given by terms corresponding to $s=3r+2$, $0\leqslant r\leqslant m$. Further,
$g_{n}(t)=\frac{1}{t^{2\mathstrut}+n^{2\mathstrut}}+\frac{2(t^{2}-n^{2}/2)}{t^{4\mathstrut}-(tn)^{2\mathstrut}+n^{4\mathstrut}}=\frac{3t^{4}}{t^{6\mathstrut}+n^{6\mathstrut}}.$
Thus we find
$\alpha(n)=-\frac{\pi}{3\sqrt{3}}-\frac{1}{6n}+\frac{1}{2}\sum\limits_{r=0}^{m}\frac{B_{6r+4}}{3r+2}\,\frac{1}{n^{6r+4}}+\frac{2(-1)^{m}}{n^{6m+4}}\int_{0}^{+\infty}\frac{t^{6m+9}}{e^{2\pi t\mathstrut}-1}\,\frac{dt}{t^{6\mathstrut}+n^{6\mathstrut}}.$
Substituting this to the expression for $\zeta^{2}(3)$ given by Corollary 1 we obtain:
$\zeta^{2}(3)=\frac{2\pi}{\sqrt{3}}\,\zeta(5)-\frac{2}{3}\,\zeta(6)+\mathfrak{S}+\\ +\sum\limits_{n=1}^{+\infty}\frac{1}{n^{5}}\biggl(-\frac{2\pi}{3\sqrt{3}}-\frac{1}{3n}+\sum\limits_{r=0}^{m}\frac{B_{6r+4}}{3r+2}\,\frac{1}{n^{6r+4}}+\frac{4(-1)^{m}}{n^{6m+4}}\int_{0}^{+\infty}\frac{t^{6m+9}}{e^{2\pi t\mathstrut}-1}\,\frac{dt}{t^{6\mathstrut}+n^{6\mathstrut}}\biggr)=\\ =\frac{2\pi}{\sqrt{3}}\,\zeta(5)-\frac{2}{3}\,\zeta(6)+\mathfrak{S}-\frac{2\pi}{3\sqrt{3}}\,\zeta(5)-\frac{1}{3}\,\zeta(6)+\sum\limits_{r=0}^{m}\frac{B_{6r+4}}{3r+2}\,\zeta(6r+9)+\\ +(-1)^{m}\int_{0}^{+\infty}\frac{4}{e^{2\pi t\mathstrut}-1}\sum\limits_{n=1}^{+\infty}\frac{(t/n)^{6}}{t^{6\mathstrut}+n^{6\mathstrut}}\,dt=\\ =\frac{4\pi}{3\sqrt{3}}\,\zeta(5)-\zeta(6)+\sum\limits_{r=0}^{m}\frac{B_{6r+4}}{3r+2}\,\zeta(6r+9)+(-1)^{m}\int_{0}^{+\infty}\frac{F_{m}(t)\,dt}{e^{2\pi t\mathstrut}-1}.$
Corollary is proved. $\square$
Taking $m=0,1,2$, we obtain the following identities:
$\zeta^{2}(3)+\zeta(6)=$ $\frac{4\pi}{3\sqrt{3}}\,\zeta(5)-\frac{1}{60}\,\zeta(9)+\mathfrak{S}+\int_{0}^{+\infty}\frac{F_{0}(t)\,dt}{e^{2\pi t\mathstrut}-1},$
$F_{0}(t)=4\sum\limits_{n=1}^{+\infty}\biggl(\frac{t}{n}\biggr)^{9}\,\frac{1}{t^{6\mathstrut}+n^{6\mathstrut}},$
$\zeta^{2}(3)+\zeta(6)=$ $\frac{4\pi}{3\sqrt{3}}\,\zeta(5)-\frac{1}{60}\,\zeta(9)+\frac{1}{66}\,\zeta(15)+\mathfrak{S}-\int_{0}^{+\infty}\frac{F_{1}(t)\,dt}{e^{2\pi t\mathstrut}-1},$
$F_{1}(t)=4\sum\limits_{n=1}^{+\infty}\biggl(\frac{t}{n}\biggr)^{15}\,\frac{1}{t^{6\mathstrut}+n^{6\mathstrut}},$
$\zeta^{2}(3)+\zeta(6)=$ $\frac{4\pi}{3\sqrt{3}}\,\zeta(5)-\frac{1}{60}\,\zeta(9)+\frac{1}{66}\,\zeta(15)-\frac{3617}{4080}\,\zeta(21)+\mathfrak{S}-\int_{0}^{+\infty}\frac{F_{2}(t)\,dt}{e^{2\pi t\mathstrut}-1},$
$F_{2}(t)=4\sum\limits_{n=1}^{+\infty}\biggl(\frac{t}{n}\biggr)^{21}\,\frac{1}{t^{6\mathstrut}+n^{6\mathstrut}}.$
[…]
We begin with one general assertion.
Theorem 4. Let $g$ be arithmetical function and let $f$ be its Dirichlet’s convolution, that is,
$f(n)=\sum\limits_{d|n}g(d),\quad\textit{for any }n\geqslant 1.$
Suppose also that $f(n),g(n)\ll_{\alpha}n^{\alpha}$ for some fixed $\alpha$, $0<\alpha<1$. Finally, for $\operatorname{Re}{z}>1+\alpha$, let
$L(z;f)=\sum\limits_{n=1}^{+\infty}\frac{f(n)}{n^{z}}.$
Then, for any $k\geqslant 1$, the following identity holds:
$L^{2}(2k;f)=\sum\limits_{m=1}^{+\infty}\frac{g(m)}{m}\biggl(\,\sum\limits_{n=1}^{+\infty}\frac{f(n)}{n^{4k-1\mathstrut}}\,a_{k}\Bigl(\frac{n}{m}\Bigr)-mL(4k;f)\biggr).$
Proof. By the Corollary 1 of Lemma 3, for any $n,q\geqslant 1$ one has
$\frac{n^{2k-1}}{q^{2k\mathstrut}+n^{2k\mathstrut}}=\frac{1}{2k}\sum\limits_{r=0}^{k-1}\varepsilon_{r}\biggl(\frac{1}{q+n\varepsilon_{r}}-\frac{1}{q-n\varepsilon_{r}}\biggr)=\frac{1}{2k}\sum\limits_{r=0}^{k-1}\biggl(\frac{1}{n+q\overline{\varepsilon}_{r}}+\frac{1}{n-q\overline{\varepsilon}_{r}}\biggr)=\\ =\frac{1}{2k}\sum\limits_{r=0}^{k-1}\biggl(\frac{1}{n+q\varepsilon_{r}}+\frac{1}{n-q\varepsilon_{r}}\biggr).$
Multiplying both parts of this identity by $f(q)$ and summing over $q\geqslant 1$ we get
$\sum\limits_{q=1}^{+\infty}\frac{n^{2k-1}f(q)}{q^{2k\mathstrut}+n^{2k\mathstrut}}=\frac{1}{2k}\sum\limits_{q=1}^{+\infty}f(q)\sum\limits_{r=0}^{k-1}\biggl(\frac{1}{n+q\varepsilon_{r}}+\frac{1}{n-q\varepsilon_{r}}\biggr).$ (14)
Since
$\biggl|\frac{1}{n+q\varepsilon_{r}}+\frac{1}{n-q\varepsilon_{r}}\biggr|=\frac{2n}{|n^{2\mathstrut}-(q\varepsilon_{r})^{2\mathstrut}|}\leqslant\biggl(\sin{\frac{\pi}{2k}}\biggr)^{-1}\frac{2n}{n^{2}+q^{2}},$ (15)
then the series
$\sum\limits_{q=1}^{+\infty}f(q)\biggl(\frac{1}{n+q\varepsilon_{r}}+\frac{1}{n-q\varepsilon_{r}}\biggr)$
converges absolutely for any $n\geqslant 1$, and for any $r$, $0\leqslant r\leqslant k-1$. Hence, Lemma 2 implies that
$\sum\limits_{q=1}^{+\infty}\frac{n^{2k-1}f(q)}{q^{2k\mathstrut}+n^{2k\mathstrut}}=\frac{1}{2k}\sum\limits_{r=0}^{k-1}\sum\limits_{q=1}^{+\infty}f(q)\biggl(\frac{1}{n+q\varepsilon_{r}}+\frac{1}{n-q\varepsilon_{r}}\biggr).$ (16)
Now let us choose a sufficiently large $M$ and split each sum over $q$ into parts corresponding to the ranges $1\leqslant q\leqslant M$ and $q>M$. The contribution given by $q>M$ is estimated by
$\sum\limits_{q>M}\frac{2n|f(q)|}{n^{2}+q^{2}}\,\biggl(\sin{\frac{\pi}{2k}}\biggr)^{-1}\ll_{k}n\sum\limits_{q>M}\frac{|f(q)|}{q^{2}}\ll_{k,\alpha}\frac{n}{M^{1-\alpha}}.$
Next, the sum over $1\leqslant q\leqslant M$ is transformed as follows:
$\sum\limits_{1\leqslant q\leqslant M}\biggl(\,\sum\limits_{m|q}g(m)\biggr)\biggl(\frac{1}{n+q\varepsilon_{r}}+\frac{1}{n-q\varepsilon_{r}}\biggr)=\\ =\sum\limits_{1\leqslant\ell m\leqslant M}g(m)\biggl(\frac{1}{n+\ell m\varepsilon_{r}}+\frac{1}{n-\ell m\varepsilon_{r}}\biggr)=\\ =\sum\limits_{1\leqslant m\leqslant M}g(m)\sum\limits_{1\leqslant\ell\leqslant M/m}\biggl(\frac{1}{n+\ell m\varepsilon_{r}}+\frac{1}{n-\ell m\varepsilon_{r}}\biggr).$ (17)
It follows from (15) that
$\sum\limits_{\ell>M/m}\biggl(\frac{1}{n+\ell m\varepsilon_{r}}+\frac{1}{n-\ell m\varepsilon_{r}}\biggr)\ll_{k}\sum\limits_{\ell>M/m}\frac{n}{n^{2\mathstrut}+(\ell m)^{2\mathstrut}}\ll_{k}\frac{n}{m^{2\mathstrut}}\sum\limits_{\ell>M/m}\frac{1}{\ell^{2}}\ll_{k}\frac{n}{Mm}.$
Therefore, the error given by the replacement the sum over $\ell$ in (17) by infinite one does not exceed in order
$\frac{n}{M}\sum\limits_{m\leqslant M}\frac{|g(m)|}{m}\ll_{k,\alpha}\frac{n}{M}\sum\limits_{m\leqslant M}\frac{1}{m^{1-\alpha}}\ll_{k,\alpha}\frac{n}{M^{1-\alpha}}$
Thus, the sum (17) coincides with
$\sum\limits_{m\leqslant M}g(m)\sum\limits_{\ell=1}^{+\infty}\biggl(\frac{1}{n+\ell m\varepsilon_{r}}+\frac{1}{n-\ell m\varepsilon_{r}}\biggr)+O\biggl(\frac{n}{M^{1-\alpha}}\biggr).$ (18)
In view of (15), for any $m$, the sum over $\ell$ does not exceed
$\biggl(\sin{\frac{\pi}{2k}}\biggr)^{-1}\sum\limits_{\ell=1}^{+\infty}\frac{2n}{n^{2}+(\ell m)^{2}}\ll_{k}\sum\limits_{\ell=1}^{+\infty}\frac{n}{n^{2\mathstrut}+(\ell m)^{2\mathstrut}}\ll_{k}\frac{n}{m^{2\mathstrut}}.$
Hence, the error that occurs when replacing the sum over $m$ in (18) with an infinite one is estimated as
$n\sum\limits_{m>M}\frac{|g(m)|}{m^{2}}\ll_{k,\alpha}\frac{n}{M^{1-\alpha}}.$
Therefore, the right-hand side of (17) coincides with
$\sum\limits_{m=1}^{+\infty}g(m)\sum\limits_{\ell=1}^{+\infty}\biggl(\frac{1}{n+\ell m\varepsilon_{r}}+\frac{1}{n-\ell m\varepsilon_{r}}\biggr)+O\biggl(\frac{n}{M^{1-\alpha}}\biggr).$
As $M\to+\infty$ tends to infinity, we obtain
$\sum\limits_{q=1}^{+\infty}f(q)\biggl(\frac{1}{n+q\varepsilon_{r}}+\frac{1}{n-q\varepsilon_{r}}\biggr)=\sum\limits_{m=1}^{+\infty}g(m)\sum\limits_{\ell=1}^{+\infty}\biggl(\frac{1}{n+\ell m\varepsilon_{r}}+\frac{1}{n-\ell m\varepsilon_{r}}\biggr).$ (19)
By Lemma 7, for any $m,n,r$, the sum over $\ell$ in (19) reduces to
$\frac{1}{m\varepsilon_{r}}\sum\limits_{\ell=1}^{+\infty}\biggl(\frac{1}{\ell+n\overline{\varepsilon}_{r}/m}-\frac{1}{\ell-n\overline{\varepsilon}_{r}/m}\biggr)=\frac{\overline{\varepsilon}_{r}}{m}\biggl(\pi\operatorname{ctg}{\frac{\pi\overline{\varepsilon}_{r}n}{m}}-\frac{m}{n\overline{\varepsilon}_{r}}\biggr)=\\ =\frac{1}{m}\biggl(\pi\overline{\varepsilon}_{r}\operatorname{ctg}{\frac{\pi\overline{\varepsilon}_{r}n}{m}}-\frac{m}{n}\biggr).$
In view of Lemma 8, for any $m,n,r$ this quantity is estimated as
$\ll\biggl(\sin{\frac{\pi}{2k+1}}\biggr)^{-1}\frac{1}{m}\,\min{\biggl(1,\frac{n}{m}\biggr)}\ll_{k}\frac{\min{(m,n)}}{m^{2}}.$
Thus, the series
$\sum\limits_{m=1}^{+\infty}g(m)\sum\limits_{\ell=1}^{+\infty}\biggl(\frac{1}{n+\ell m\varepsilon_{r}}+\frac{1}{n-\ell m\varepsilon_{r}}\biggr)=\sum\limits_{m=1}^{+\infty}\frac{g(m)}{m}\biggl(\pi\overline{\varepsilon}_{r}\operatorname{ctg}{\frac{\pi\overline{\varepsilon}_{r}n}{m}}-\frac{m}{n}\biggr)$ (20)
converges absolutely. Using Lemma 2 and changing the order of summation, from (16), (19) and (20) we get
$\sum\limits_{q=1}^{+\infty}\frac{n^{2k-1}f(q)}{q^{2k\mathstrut}+n^{2k\mathstrut}}=\frac{1}{2}\sum\limits_{m=1}^{+\infty}\frac{g(m)}{m}\,\frac{1}{k}\sum\limits_{r=0}^{k-1}\biggl(\pi\overline{\varepsilon}_{r}\operatorname{ctg}{\frac{\pi\overline{\varepsilon}_{r}n}{m}}-\frac{m}{n}\biggr).$ (21)
Since
$\biggl|\pi\overline{\varepsilon}_{r}\operatorname{ctg}{\frac{\pi\overline{\varepsilon}_{r}n}{m}}-\frac{m}{n}\biggr|\ll_{k}\min{\biggl(1,\frac{n}{m}\biggr)}\ll_{k}\frac{\min{(m,n)}}{m},$
then
$\frac{1}{k}\sum\limits_{r=0}^{k-1}\biggl(\pi\overline{\varepsilon}_{r}\operatorname{ctg}{\frac{\pi\overline{\varepsilon}_{r}n}{m}}-\frac{m}{n}\biggr)=\overline{a}_{k}\biggl(\frac{n}{m}\biggr)-\frac{m}{n}=a_{k}\biggl(\frac{n}{m}\biggr)-\frac{m}{n}\ll\frac{\min{(m,n)}}{m}$ (22)
and hence
$\biggl|a_{k}\biggl(\frac{n}{m}\biggr)\biggr|\ll\frac{m}{n}+\frac{\min{(m,n)}}{m}.$ (23)
Multiplying both parts of (21) by $f(n)\cdot n^{-(4k-1)}$ and summing over $n\geqslant 1$, we find
$\sum\limits_{n=1}^{+\infty}\frac{f(n)}{n^{2k}}\sum\limits_{q=1}^{+\infty}\frac{f(q)}{q^{2k\mathstrut}+n^{2k\mathstrut}}=\frac{1}{2}\sum\limits_{n=1}^{+\infty}\frac{f(n)}{n^{4k-1}}\sum\limits_{m=1}^{+\infty}\frac{g(m)}{m}\biggl(a_{k}\biggl(\frac{n}{m}\biggr)-\frac{m}{n}\biggr).$ (24)
By Lemma 1, the left hand side of (24) is equal to
$\sum\limits_{n,q=1}^{+\infty}\frac{f(n)f(q)}{(nq)^{2k\mathstrut}}\cdot\frac{q^{2k}}{q^{2k\mathstrut}+n^{2k\mathstrut}}=\frac{1}{2}\biggl(\,\sum\limits_{q=1}^{+\infty}\frac{f(q)}{q^{2k\mathstrut}}\biggr)^{2}=\frac{1}{2}\,L^{2}(2k;f).$
In view of (22), the double series in the right-hand side of (24) converges absolutely. Using Lemma 2, we get
$\frac{1}{2}\sum\limits_{n=1}^{+\infty}\frac{f(n)}{n^{4k-1}}\sum\limits_{m=1}^{+\infty}\frac{g(m)}{m}\biggl(a_{k}\biggl(\frac{n}{m}\biggr)-\frac{m}{n}\biggr)=\frac{1}{2}\sum\limits_{m=1}^{+\infty}\frac{g(m)}{m}\sum\limits_{n=1}^{+\infty}\frac{f(n)}{n^{4k-1}}\biggl(a_{k}\biggl(\frac{n}{m}\biggr)-\frac{m}{n}\biggr).$
Next, in view of (23), the inner sum splits into two absolutely convergent series. Thus we obtain:
$\frac{1}{2}\,L^{2}(2k;f)=\frac{1}{2}\sum\limits_{m=1}^{+\infty}\frac{g(m)}{m}\biggl(\,\sum\limits_{n=1}^{+\infty}\frac{f(n)}{n^{4k-1}}\,a_{k}\biggl(\frac{n}{m}\biggr)-mL(4k;f)\biggr).$
Theorem is proved. $\square$
Corollary 1. Under the assumptions of Theorem 4, the following identity holds:
$L^{2}(2;f)=\sum\limits_{m=1}^{+\infty}\frac{g(m)}{m}\biggl(\,2\pi\sum\limits_{n=1}^{+\infty}\frac{f(n)}{n^{3\mathstrut}\bigl(e^{2\pi n/m}-1\bigr)}-mL(4;f)+\pi L(3;f)\biggr).$
Proof. Indeed, taking $k=1$ we get
$L^{2}(2;f)=\sum\limits_{m=1}^{+\infty}\frac{g(m)}{m}\biggl(\,\sum\limits_{n=1}^{+\infty}\frac{f(n)}{n^{3\mathstrut}}\,a_{1}\biggl(\frac{n}{m}\biggr)-mL(4;f)\biggr).$
Now it remains to note that
$a_{1}(w)=\pi i\operatorname{ctg}{(\pi iw)}=\pi\operatorname{cth}{(\pi w)}=\pi\biggl(1+\frac{2}{e^{2\pi w}-1}\biggr).\quad\square$
In the below, we give some particular cases of Corollary 1.
$1^{\circ}$. Taking $g(m)=\mu(m)$, we get
$f(n)=\begin{cases}1,&\text{ if }n=1,\\ 0,&\text{ if }n>1.\end{cases}$
Hence, $L(z;f)\equiv 1$ for any $z$. Therefore, we have
$\sum\limits_{m=1}^{+\infty}\frac{\mu(m)}{m}\biggl(\frac{2\pi}{e^{2\pi/m}-1}+\pi-m\biggr)=1.$
If we express the left part as the limit of partial sums over $1\leqslant m\leqslant M$ when $M\to+\infty$ and use the fact that
$\lim_{M\to+\infty}\sum\limits_{1\leqslant m\leqslant M}\frac{\mu(m)}{m}=0,$
we find
$\sum\limits_{m=1}^{+\infty}\mu(m)\biggl(\frac{2\pi/m}{e^{2\pi/m}-1}\;-\;1\biggr)=1.$
$2^{\circ}$. Let $\nu\geqslant 1$ be integer. Setting $g(m)=\tau_{\nu}(m)$, we get $f(n)=\tau_{\nu+1}(n)$ and $L(z;f)=\zeta^{\nu+1}(z)$. Hence
$\zeta^{2\nu+2}(2)=\sum\limits_{m=1}^{+\infty}\frac{\tau_{\nu}(m)}{m}\biggl(\,\sum\limits_{n=1}^{+\infty}\frac{\tau_{\nu+1}(n)}{n^{3\mathstrut}}\,\frac{2\pi}{e^{2\pi n/m}-1}-m\zeta^{\nu+1}(4)+\pi\zeta^{\nu+1}(3)\biggr).$
In particular, for $\nu=1$ and $\nu=2$ we correspondingly have
$\zeta^{4}(2)=\sum\limits_{m=1}^{+\infty}\frac{1}{m}\biggl(\,\sum\limits_{n=1}^{+\infty}\frac{\tau(n)}{n^{3\mathstrut}}\,\frac{2\pi}{e^{2\pi n/m}-1}-m\zeta^{2}(4)+\pi\zeta^{2}(3)\biggr),$
$\zeta^{6}(2)=\sum\limits_{m=1}^{+\infty}\frac{\tau(m)}{m}\biggl(\,\sum\limits_{n=1}^{+\infty}\frac{\tau_{3}(n)}{n^{3\mathstrut}}\,\frac{2\pi}{e^{2\pi n/m}-1}-m\zeta^{3}(4)+\pi\zeta^{3}(3)\biggr).$
$3^{\circ}$. Setting $g(m)=\mu^{2}(m)$, we get
$f(n)=\sum\limits_{d|n}\mu^{2}(d)=2^{\omega(n)},\quad L(z;f)=\frac{\zeta^{2}(z)}{\zeta(2z)}\quad\text{for}\quad\operatorname{Re}z>1,$
and hence
$\frac{\zeta^{2}(2)}{\zeta(4)}=\sum\limits_{m=1}^{+\infty}\frac{\mu^{2}(m)}{m}\biggl(\,\sum\limits_{n=1}^{+\infty}\frac{2^{\omega(n)}}{n^{3\mathstrut}}\,\frac{2\pi}{e^{2\pi n/m}-1}-m\,\frac{\zeta^{2}(4)}{\zeta(8)}+\pi\,\frac{\zeta^{2}(3)}{\zeta(6)}\biggr).$
$4^{\circ}$. Setting
[…]
Thus
$f(n)=\mu^{2}(n),\quad L(z;f)=\frac{\zeta(z)}{\zeta(2z)},\quad\text{for}\quad\operatorname{Re}z>1,$
and hence
$\biggl(\frac{\zeta(2)}{\zeta(4)}\biggr)^{2}=\sum\limits_{m=1}^{+\infty}\frac{\mu(m)}{m^{2\mathstrut}}\biggl(\,\sum\limits_{n=1}^{+\infty}\frac{\mu^{2}(n)}{n^{3\mathstrut}}\,\frac{2\pi}{e^{2\pi n/m^{2}}-1}-m^{2}\,\frac{\zeta(4)}{\zeta(8)}+\pi\,\frac{\zeta(3)}{\zeta(6)}\biggr).$
$5^{\circ}$. Setting $g(m)=\tau(m^{2})$ and using the identities
$\frac{\zeta^{3}(z)}{\zeta(2z)}=\sum\limits_{m=1}^{+\infty}\frac{\tau(m^{2})}{m^{z}},\quad\frac{\zeta^{4}(z)}{\zeta(2z)}=\sum\limits_{m=1}^{+\infty}\frac{\tau^{2}(m)}{m^{z}},\quad\operatorname{Re}z>1,$
we conclude that
$f(n)=\tau^{2}(n),\quad L(z;f)=\frac{\zeta^{4}(z)}{\zeta(2z)}\quad\text{for}\quad\operatorname{Re}z>1$
and hence
$\frac{\zeta^{8}(2)}{\zeta^{2}(4)}=\sum\limits_{m=1}^{+\infty}\frac{\tau(m^{2})}{m}\biggl(\,\sum\limits_{n=1}^{+\infty}\frac{\tau^{2}(n)}{n^{3\mathstrut}}\,\frac{2\pi}{e^{2\pi n/m}-1}-m\,\frac{\zeta^{4}(4)}{\zeta(8)}+\pi\,\frac{\zeta^{4}(3)}{\zeta(6)}\biggr).$
$6^{\circ}$. Setting $g(m)=\lambda(m)$, we get
$f(n)=\begin{cases}1,&\text{if }n\text{ is a perfect square},\\ 0,&\text{otherwise},\end{cases}\quad L(z;f)=\zeta(2z)\quad\text{for}\quad\operatorname{Re}z>\frac{1}{2},$
and hence
$\zeta^{2}(4)=\sum\limits_{m=1}^{+\infty}\frac{\lambda(m)}{m}\biggl(\,\sum\limits_{n=1}^{+\infty}\frac{1}{n^{6\mathstrut}}\,\frac{2\pi}{e^{2\pi n^{2\mathstrut}/m}-1}-m\,\zeta(8)+\pi\,\zeta(6)\biggr).$
$7^{\circ}$. Setting $g(m)=\mu(m)/m$, we get
$f(n)=\frac{\varphi(n)}{n},\quad L(z;f)=\frac{\zeta(z)}{\zeta(z+1)}\quad\text{for}\quad\operatorname{Re}z>1$
and thus
$\biggl(\frac{\zeta(2)}{\zeta(3)}\biggr)^{2}=\sum\limits_{m=1}^{+\infty}\frac{\mu(m)}{m^{2\mathstrut}}\biggl(\,\sum\limits_{n=1}^{+\infty}\frac{\varphi(n)}{n^{4\mathstrut}}\,\frac{2\pi}{e^{2\pi n/m}-1}-m\,\frac{\zeta(4)}{\zeta(5)}+\pi\,\frac{\zeta(3)}{\zeta(4)}\biggr).$
$8^{\circ}$. Setting $g(m)=\Lambda(m)$, we get $f(n)=\ln{n}$, $L(z;f)=-\zeta^{\prime}(z)$, so we have
$\bigl(\zeta^{\prime}(2)\bigr)^{2}=\sum\limits_{m=1}^{+\infty}\frac{\Lambda(m)}{m}\biggl(\,\sum\limits_{n=1}^{+\infty}\frac{\ln{n}}{n^{3\mathstrut}}\,\frac{2\pi}{e^{2\pi n/m}-1}+m\,\zeta^{\prime}(4)-\pi\,\zeta^{\prime}(3)\biggr).$
$9^{\circ}$. Suppose that $k\geqslant 1$ and let $\Lambda_{k}$ be the generalized von Mangoldt function defined by the following relation:
$\sum\limits_{m=1}^{+\infty}\frac{\Lambda_{k}(m)}{m^{z}}=(-1)^{k}\,\frac{\zeta^{(k)}(z)}{\zeta(z)}.$
Taking $g(m)=\Lambda_{k}(m)$ we get $f(n)=(-1)^{k}(\ln{n})^{k}$, $L(z;f)=(-1)^{k}\zeta^{(k)}(z)$ and therefore
$\bigl(\zeta^{(k)}(2)\bigr)^{2}=(-1)^{k}\sum\limits_{m=1}^{+\infty}\frac{\Lambda_{k}(m)}{m}\biggl(\,\sum\limits_{n=1}^{+\infty}\frac{(\ln{n})^{k}}{n^{3\mathstrut}}\,\frac{2\pi}{e^{2\pi n/m}-1}-m\,\zeta^{(k)}(4)+\pi\,\zeta^{(k)}(3)\biggr).$
$10^{\circ}$. Let $\chi_{4}$ be non-principal Dirichlet character modulo $4$, that is
[…]
Then, setting $g(m)=\chi_{4}(m)$, we get $f(n)=r(n)/4$, where $r(n)$ is the number of representations of $n$ by the sum of two squares of integers. Then
$L(z;f)=\frac{1}{4}\,\zeta(z)L(z),\quad\text{where}\quad L(z)=L(z;\chi_{4})=\sum\limits_{m=1}^{+\infty}\frac{\chi_{4}(m)}{m^{z}}$
is Dirichlet $L$-function corresponding to $\chi_{4}$. After easy calculations, we get
$\zeta^{2}(2)\beta^{2}=4\sum\limits_{m=1}^{+\infty}\frac{\chi_{4}(m)}{m}\biggl(\,\sum\limits_{n=1}^{+\infty}\frac{r(n)}{n^{3\mathstrut}}\,\frac{2\pi}{e^{2\pi n/m}-1}-m\,\zeta(4)L(4)+\pi\,\zeta(3)L(3)\biggr),$
where
$\beta=L(2)=\sum\limits_{n=0}^{+\infty}\frac{(-1)^{n}}{(2n+1)^{2}}\quad\text{is Catalan constant.}$
$11^{\circ}$. Let $g(m)=c_{m}(a)$, where
$c_{m}(a)=\sum\limits_{\begin{subarray}{c}\nu=1\\ (\nu,m)=1\end{subarray}}^{m}e^{2\pi i\,\frac{\scriptstyle\nu a}{\scriptstyle m}}\quad\text{is Ramanujan's sum}$
(see [8]; here and below we suppose that $a\geqslant 1$ is a fixed integer). Then
$f(n)=\sum\limits_{m|n}g(m)=\sum\limits_{m|n}c_{m}(a)=\begin{cases}n,&\text{if }n|a,\\ 0,&\text{otherwise}.\end{cases}$
Hence,
$L(z;f)=\sum\limits_{n|a}\frac{n}{n^{z}}=\sigma_{1-z}(a)=\frac{\sigma_{z-1}(a)}{a^{z-1}}$
and, in particular,
$L(2;f)=\frac{\sigma(a)}{a},\quad L(3;f)=\frac{\sigma_{2}(a)}{a^{2\mathstrut}},\quad L(4;f)=\frac{\sigma_{3}(a)}{a^{3\mathstrut}}.$
Thus we find
$\biggl(\frac{\sigma(a)}{a}\biggr)^{2}=\sum\limits_{m=1}^{+\infty}\frac{c_{m}(a)}{m}\biggl(\,\sum\limits_{d|a}\frac{2\pi}{d^{2}(e^{2\pi d/m}-1)}-m\,\frac{\sigma_{3}(a)}{a^{3\mathstrut}}+\pi\,\frac{\sigma_{2}(a)}{a^{2\mathstrut}}\biggr).$
Note that
$\sum\limits_{m=1}^{+\infty}\frac{c_{m}(a)}{m}=0\quad\text{for any }a\geqslant 1.$
Thus we find
$\sigma^{2}(a)=a^{2}\sum\limits_{m=1}^{+\infty}c_{m}(a)\biggl(\,\sum\limits_{d|a}\frac{1}{d^{2\mathstrut}}\,\frac{2\pi/m}{e^{2\pi d/m}-1}-\frac{\sigma_{3}(a)}{a^{3\mathstrut}}\biggr).$
In the case $a=1$, using the fact $c_{m}(1)=\mu(m)$, we get Cauchy-Lerch-Ramanujan identity from $1^{\circ}$.
Theorem 5. Under the assumptions of Theorem 4, for any $k\geqslant 2$, the following identity holds
$L^{2}(2k-1;f)=\sum\limits_{m=1}^{+\infty}\frac{g(m)}{m}\sum\limits_{n=1}^{+\infty}\frac{f(n)}{n^{4k-3}}\,b_{k}\biggl(\frac{n}{m}\biggr).$
Proof. Taking $\ell=1$ in Corollary 2 of Lemma 3, we get
$\frac{qn^{2k-2}}{q^{2k\mathstrut}+n^{2k\mathstrut}}=\frac{1}{2k}\sum\limits_{r=0}^{2k-1}\varepsilon_{r}^{2}\biggl(\frac{1}{q}-\frac{1}{q+n\varepsilon_{r}}\biggr)$
and therefore
$\sum\limits_{q=1}^{+\infty}\frac{n^{2k-2}\,qf(q)}{q^{2k\mathstrut}+n^{2k\mathstrut}}=\frac{1}{2k}\sum\limits_{q=1}^{+\infty}f(q)\sum\limits_{r=0}^{2k-1}\varepsilon_{r}^{2}\biggl(\frac{1}{q}-\frac{1}{q+n\varepsilon_{r}}\biggr).$
Using the inequality
$|q+n\varepsilon_{r}|\geqslant(q+n)\sin{\frac{\pi}{4k}}$
and follow word-by-word to the proof of Theorem 4, we obtain
$\sum\limits_{q=1}^{+\infty}\frac{n^{2k-2}\,qf(q)}{q^{2k\mathstrut}+n^{2k\mathstrut}}=\frac{1}{2k}\sum\limits_{r=0}^{2k-1}\varepsilon_{r}^{2}\sum\limits_{m,\ell=1}^{+\infty}g(m)\biggl(\frac{1}{m\ell}-\frac{1}{m\ell+n\varepsilon_{r}}\biggr)=\\ =\frac{1}{2k}\sum\limits_{r=0}^{2k-1}\varepsilon_{r}^{2}\sum\limits_{m=1}^{+\infty}\frac{g(m)}{m}\sum\limits_{\ell=1}^{+\infty}\biggl(\frac{1}{\ell}-\frac{1}{\ell+n\varepsilon_{r}/m}\biggr)=\frac{1}{2k}\sum\limits_{r=0}^{2k-1}\varepsilon_{r}^{2}\sum\limits_{m=1}^{+\infty}\frac{g(m)}{m}\biggl(\psi\biggl(\frac{n\varepsilon_{r}}{m}\biggr)+\gamma+\frac{m}{n\varepsilon_{r}}\biggr)=\\ =\frac{1}{2}\sum\limits_{m=1}^{+\infty}\frac{g(m)}{m}\,\frac{1}{k}\sum\limits_{r=0}^{2k-1}\varepsilon_{r}^{2}\biggl(\psi\biggl(\frac{n\varepsilon_{r}}{m}\biggr)+\gamma+\frac{m}{n\varepsilon_{r}}\biggr)=\frac{1}{2}\sum\limits_{m=1}^{+\infty}\frac{g(m)}{m}\,b_{k}\biggl(\frac{n}{m}\biggr).$
Multiplying both parts by $f(n)\cdot n^{-(4k-3)}$ and summing over $n$ we find
$\frac{1}{2}\sum\limits_{n=1}^{+\infty}\frac{f(n)}{n^{4k-3}}\sum\limits_{m=1}^{+\infty}\frac{g(m)}{m}\,b_{k}\biggl(\frac{n}{m}\biggr)=\frac{1}{2}\sum\limits_{m=1}^{+\infty}\frac{g(m)}{m}\sum\limits_{n=1}^{+\infty}\frac{f(n)}{n^{4k-3}}\,b_{k}\biggl(\frac{n}{m}\biggr)=\\ \frac{1}{2}\sum\limits_{n=1}^{+\infty}\frac{f(n)}{n^{4k-3}}\sum\limits_{q=1}^{+\infty}\frac{n^{2k-2}\,qf(q)}{q^{2k\mathstrut}+n^{2k\mathstrut}}=\sum\limits_{q,n=1}^{+\infty}\frac{f(q)f(n)}{(qn)^{2k-1}}\,\frac{q^{2k}}{q^{2k\mathstrut}+n^{2k\mathstrut}}=\\ =\frac{1}{2}\biggl(\,\sum\limits_{n=1}^{+\infty}\frac{f(n)}{n^{2k-1}}\biggr)^{2}=\frac{1}{2}\,L^{2}(2k-1;f).\quad\square$
Corollary. Let $\varepsilon=e^{\pi i/4}$, $\omega=e^{\pi i/6}$. Then, under the assumptions of Theorem 4, the following identities hold:
$L^{2}(3;f)=$ $\frac{i}{2}\sum\limits_{m=1}^{+\infty}\frac{g(m)}{m}\sum\limits_{n=1}^{+\infty}\frac{f(n)}{n^{5}}\,\biggl(\psi\Bigl(\frac{n\varepsilon}{m}\Bigr)+\psi\Bigl(-\frac{n\varepsilon}{m}\Bigr)-\psi\Bigl(\frac{n\overline{\varepsilon}}{m}\Bigr)-\psi\Bigl(-\frac{n\overline{\varepsilon}}{m}\Bigr)\biggr),$
$L^{2}(5;f)=$ $\frac{1}{3}\sum\limits_{m=1}^{+\infty}\frac{g(m)}{m}\sum\limits_{n=1}^{+\infty}\frac{f(n)}{n^{9}}\,\biggl\{\omega^{2}\biggl(\psi\Bigl(\frac{n\omega}{m}\Bigr)+\psi\Bigl(-\frac{n\omega}{m}\Bigr)\biggr)+\overline{\omega}^{2}\biggl(\psi\Bigl(\frac{n\overline{\omega}}{m}\Bigr)+\psi\Bigl(-\frac{n\overline{\omega}}{m}\Bigr)\biggr)$-
- $\psi\Bigl(\frac{ni}{m}\Bigr)-\psi\Bigl(-\frac{ni}{m}\Bigr)\biggr\}.$
Theorem 6. Under the assumptions of Theorem 4, for any $k\geqslant 1$, the following identity holds:
$\frac{1}{2}\,L^{2}(2k+1;f)=\sum\limits_{m=1}^{+\infty}\frac{g(m)}{m}\biggl(\,\sum\limits_{n=1}^{+\infty}\frac{f(n)}{n^{4k+1}}\,c_{k}\Bigl(\frac{n}{m}\Bigr)-L(4k+2;f)\biggr).$
Proof is based on the expansion
$\frac{n^{2k}}{q^{2k+1\mathstrut}+n^{2k+1\mathstrut}}=\frac{1}{2k+1}\sum\limits_{r=0}^{2k}\omega_{r}\biggl(\frac{1}{q}-\frac{1}{q-n\omega_{r}}\biggr)$
(see Corollary 1 to Lemma 4) and follows word-by-word to the proofs of Theorem 4, 5.
Corollary. Let $\varpi=e^{\pi i/3}$. Then, under the assumptions of Theorem 4, the following identity holds:
$\frac{1}{2}\,L^{2}(3;f)=\sum\limits_{m=1}^{+\infty}\frac{g(m)}{m}\biggl(\frac{1}{3}\sum\limits_{n=1}^{+\infty}\frac{1}{n^{5}}\biggl\{\varpi\psi\Bigl(-\frac{n\varpi}{m}\Bigr)+\overline{\varpi}\psi\Bigl(-\frac{n\overline{\varpi}}{m}\Bigr)-\psi\Bigl(\frac{n}{m}\Bigr)\biggr\}-L(6;f)\biggr).$